OpenStudy (anonymous):

for this quadratic equation, does a=6,b=-3, and c=-2?

OpenStudy (anonymous):

\[6x ^{2}-3=-2x\]

OpenStudy (anonymous):

no....b=2, c=-3

OpenStudy (anonymous):

Or should b or c be positive?

OpenStudy (anonymous):

once you put everything together on one side, you get: 6x^2+2x-3=0

OpenStudy (anonymous):

6x^2+2x-3, a=6 b=2 c=-3

OpenStudy (anonymous):

ohh because it has x with it?

OpenStudy (anonymous):

exactly :)

OpenStudy (anonymous):

think of the generic form of polynomials like that: ax^2+bx+c

OpenStudy (anonymous):

thanks! I'm going to try and solve it, would you mind checking it to see if I have the correct answer? :)

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

\[4+ or - \sqrt{76}\over 12\] not sure if this can be simplified or not

OpenStudy (anonymous):

D = 4+72 = 76 > 0 --> 2 real number (-2 +- sqrt(76)) 4

OpenStudy (anonymous):

yup ur rite sorry im typo, denominator is 12

OpenStudy (anonymous):

that should be a -2, not a 4. remember the formula is [-b +/- sqrt(b^2-4ac)]/2a

OpenStudy (anonymous):

ohh ok, it is -2 instead of 4 though?

OpenStudy (anonymous):

what is b in this equation? its 2...so-b is -2

OpenStudy (anonymous):

in the original equation it is -2..but does it turn to positive 2 when we solve it?

OpenStudy (anonymous):

in the original is 2, coz its -b so it's -2

OpenStudy (anonymous):

ohhhh right. My mistake. thanks for clarifying!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

wait a sec..i tried to submit the answer and it says it's not in simplest form?

OpenStudy (anonymous):

76=4x19. You can pull out a 2 there and then factor a 2 from the whole numerator and then you end up with: [-1+/-sqrt19]/6

OpenStudy (anonymous):

Thanks!